# IBNorm: Information-Bottleneck Inspired Normalization for Representation Learning

### v1

@@ Abstract
Normalization is fundamental to deep learning, but existing approaches such as BatchNorm, LayerNorm, and RMSNorm are variance-centric by enforcing zero mean and unit variance, stabilizing training without controlling how representations capture task-relevant information. We propose IB-Inspired Normalization (IBNorm), a simple yet powerful family of methods grounded in the Information Bottleneck principle. IBNorm introduces bounded compression operations that encourage embeddings to preserve predictive information while suppressing nuisance variability, yielding more informative representations while retaining the stability and compatibility of standard normalization. Theoretically, we prove that IBNorm achieves a higher IB value and tighter generalization bounds than variance-centric methods. Empirically, IBNorm consistently outperforms BatchNorm, LayerNorm, and RMSNorm across large-scale language models (LLaMA, GPT-2) and vision models (ResNet, ViT), with mutual information analysis confirming superior information bottleneck behavior. Code will be released publicly.
@@ 1 Introduction
Normalization has become a cornerstone of deep learning, credited for stabilizing and accelerating training across domains. Techniques such as Batch Normalization (BN) (Ioffe & Szegedy, 2015), Layer Normalization (LN) (Ba et al., 2016), and RMSNorm (Zhang & Sennrich, 2019) are now standard in architectures spanning vision and language (Vaswani et al., 2017; Dosovitskiy et al., 2020; He et al., 2016a), underscoring normalization as a key driver of training stability and acceleration.
@@
Despite this success, existing normalization methods share a fundamental limitation: they are inherently variance-centric. By enforcing zero mean and unit variance, followed by rescaling and shifting, they improve the conditioning of the optimization problem (Huang et al., 2023; 2018; He et al., 2016b; Desjardins et al., 2015). However, this process operates purely on first- and second-order statistics, without offering any guidance on (intermediate) representation learning. Indeed, two representations may share identical mean and variance, but can encode drastically different amounts of task-relevant information. This raises a key question: can normalization be re-designed not only to stabilize training but also to shape representations toward sufficiency and generalization?
Early attempts have explored this direction. For instance, Eftekhari & Papyan (2025) encouraged Gaussian-like features via power transforms and additive noise, arguing that Gaussianity aids representation and noise acts as implicit regularization. However, such approaches lack a principled information-theoretic foundation and fail to explicitly balance preserving predictive information against suppressing nuisance factors, often yielding suboptimal embeddings for downstream tasks.
@@
Contributions. In this work, we move beyond variance-centric normalization and introduce IBNorm, a simple yet powerful normalization method grounded by the Information Bottleneck (IB) principle (Shamir et al., 2010; Tishby & Zaslavsky, 2015; Shwartz-Ziv & Tishby, 2017). Moreover, we also theoretically justify the superiority of IBNorm and its better generalization performance.
@@
First, we introduce an IB perspective to normalization, and propose IBNorm for generalization improvement. The IB principle seeks representations that preserve task-relevant information while discarding nuisances, thereby connecting normalization to generalization. Within this framework, we decompose existing normalization methods into three steps: (i) grouping features (e.g., across batch, channel, or dimension), (ii) standardization, and (iii) recovery via re-scaling and shifting. Our analysis reveals that the normalization operation itself is the key bottleneck governing information flow. Building on this insight, IBNorm (IBNorm-S, IBNorm-L, IBNorm-T) introduces compression operators that selectively compress activations toward their means, providing explicit control over sufficiency–redundancy tradeoffs while remaining drop-in compatible with existing architectures.
@@
Second, we provide theoretical guarantees. Under the IB framework, we prove that IBNorm achieves a strictly larger IB value than variance-centric methods like LN and BN, thereby better balancing predictive sufficiency with nuisance suppression. Moreover, we establish a provably tighter generalization bound, explaining why IBNorm consistently outperforms standard normalization in practice.
@@
Extensive experiments show the effectiveness of IBNorm across architectures and modalities. In language modeling, integrating IBNorm into LLaMA (Touvron et al., 2023b) and GPT-2 (Radford et al., 2019) consistently outperforms LN, RMSNorm, and NormalNorm on LLM Leaderboard I/II by up to 8.75%. In vision, applying IBNorm to ResNet (He et al., 2016a) and ViT (Dosovitskiy et al., 2020) yields substantial gains, e.g., 3.98% improvement on ImageNet (Deng et al., 2009) (ResNet-50), and up to 8.17% improvement on ImageNet (ViT).
@@ 2 Related Work
Normalization plays a central role in modern deep learning, enabling faster convergence and improved generalization across natural language processing (NLP) (Devlin et al., 2019; Vaswani et al., 2017; Touvron et al., 2023a; b; Dubey et al., 2024; Bai et al., 2023; Liu et al., 2024) and computer vision (CV) (He et al., 2016a; Dosovitskiy et al., 2020; Liu et al., 2022).
@@
Batch Normalization (BN) (Ioffe & Szegedy, 2015) normalizes activations across mini-batches to stabilize training, with numerous extensions such as Mean-only BN (Salimans & Kingma, 2016), $L^{p}$-Norm BN (Liao et al., 2016), Conditional BN (De Vries et al., 2017), and Whitening BN (Siarohin et al., 2018). Layer Normalization (LN) (Ba et al., 2016) was proposed for sequential and small-batch settings and has inspired variants including Dynamic LN (Kim et al., 2017), RMSNorm (Zhang & Sennrich, 2019), and Adaptive LN (Xu et al., 2019), which are now standard in Transformers (Vaswani et al., 2017; Dosovitskiy et al., 2020). Group Normalization (GN) (Wu & He, 2018) and Batch Group Normalization (BGN) (Summers & Dinneen, 2019) extend this idea by combining grouping across features, channels, or batches. Collectively, these methods underscore normalization as a structural tool for scaling depth and stabilizing optimization.
@@
Beyond variance-centric approaches, some works attempt to explicitly shape activation distributions. For example, Eftekhari & Papyan (2025) encourage Gaussian-like features via power transforms with additive noise, arguing that Gaussianity enhances representational capacity while noise acts as implicit regularization. However, this approach faces key limitations: it does not optimize mutual information with targets, incurs overhead in estimating power parameters, and introduces stochasticity that can destabilize training (see Appendix B).
@@
Broadly, most normalization methods rely on re-centering and rescaling to mitigate internal covariate shift, yet they overlook the information-theoretic properties of representations, such as sufficiency and redundancy. This gap motivates our work. We propose IBNorm, which augments conventional normalization with compression operation designed to regulate information flow, guiding activations toward IB-optimal representations that enhance both generalization and robustness.
@@ 3 Preliminary
Here we introduce information bottleneck (IB) framework (Tishby et al., 2000; Shamir et al., 2010; Tishby & Zaslavsky, 2015) which forms the basis of our approach in Sec. 4.
Consider two continuous random variables $X$ and $Y$ with joint probability densities $p(x,y)$ over the space $\mathcal{X}\times\mathcal{Y}$. Then, we can define the mutual information between $X$ and $Y$:
$I(X;Y)=\int_{x\in\mathcal{X}}\int_{y\in\mathcal{Y}}p(x,y)\,\log\frac{p(x,y)}{p(x)\,p(y)}\,dx\,dy.$ (1)
Mutual information measures dependencies between random variables, and can be understood as how much knowing $X$ reduces the uncertainty in $Y$ or vice versa.
Building on this, the IB framework formulates representation learning as a trade-off between two goals: 1) minimality which compresses $X$ by reducing $I(X;T)$; and 2) sufficiency which preserves information about $Y$ by maximizing $I(Y;T)$. This is captured by the IB objective (Tishby & Zaslavsky, 2015):
$\max_{T}\Big[I(Y;T)-\beta I(X;T)\Big].$ (2)
where $\beta>0$. Here, $T$ must satisfy the Markov chain TXYT-X-Y, and the optimal representation $T^{*}$ is given by the conditional distribution $p_{T^{*}|X}(t|x)$ with marginal $p_{T^{*}}(t)=\int p_{T^{*}|X}(t|x)p_{X}(x)dx$ govern by self-consistency equations (Tishby et al., 2000; Shamir et al., 2010).
Intuitively, maximizing $I(Y;T)$ ensures that the representation $T$ preserving all information relevant for predicting the target variable $Y$, thereby enhancing predictive power and interpretability. Concurrently, penalizing $I(X;T)$ enforces compression, which eliminates irrelevant variability in the representation and, in turn, enhances generalization and robustness of the representation (Kawaguchi et al., 2023; Alemi et al., 2016; Goldfeld & Polyanskiy, 2020; Saxe et al., 2019; Federici et al., 2020).
@@ 4 Methodology
Inspired by the IB principle introduced in Sec. 3, we now present how it motivates a novel and more effective normalization strategy — an essential component of modern deep learning architectures.
@@ 4.1 Normalization Design via IB Principle
Consider a data space $\mathcal{X}$ and label space $\mathcal{Y}$ with a fixed joint distribution $\mathbb{P}(X,Y)\in\mathcal{P}(\mathcal{X}\times\mathcal{Y})$. A feedforward neural network $f(\cdot;\theta)$, parameterized by $\theta$, maps an input $X\in\mathcal{X}$ to a prediction $\hat{Y}=f(X;\theta)$, aiming to approximate the ground-truth label $Y$. Here, $f(\cdot;\theta)$ can represent various architectures, such as convolutional neural networks (e.g., ResNet (He et al., 2016a)) and transformer-based networks (e.g., large language models like (Touvron et al., 2023b) or vision transformer (Dosovitskiy et al., 2020)).
For an $L$-layer network equipped with normalization, we write the intermediate representation after the $l$-th normalization as
$T_{l}:=\Phi_{l}\circ h_{l}\circ\cdots\circ\Phi_{1}\circ h_{1}(X)\in\mathcal{T}_{l},\quad l\in\{1,\cdots,L\},$ (3)
where $\mathcal{T}_{l}$ is a measurable space, $h_{i}$ is a transformation block with parameter $\theta_{i}$, and $\Phi_{i}$ is the normalization layer, e.g., layer normalization. We use symbol $\circ$ to represent the composition of functions. Following (Huang et al., 2023), any normalization layer $\Phi$ can be decomposed into three steps:
@@
This decomposition unifies existing normalization methods. For instance, BN (Ioffe & Szegedy, 2015), LN (Ba et al., 2016), and GN (Wu & He, 2018) differ mainly in their choice of NAP ($\zeta$): BN normalizes across the batch dimension, LN across the feature dimension, and GN across groups of features. In all cases, the NOP enforces zero-mean and unit-variance statistics, while the NRR rescales and shifts the result. Crucially, among them, NAP ($\zeta$) is often dictated by the network architecture or task domain (e.g., batch-based normalization favored by CNNs, feature-based normalization for Transformers). Likewise, NRR ($\eta$) is linear and invertible, and therefore only re-parameterizes activations without altering their information content. This leaves the NOP ($\psi$) as the main degree of freedom: it governs how information is filtered and thus determines whether normalization merely stabilizes optimization or also enhances the quality of learned representations.
However, most existing NOP ($\psi$) operations are fundamentally variance-centric, enforcing zero-mean and unit-variance scaling of activations. While stabilizing and accelerating training, it overlooks whether the resulting representations retain task-relevant information. From the IB perspective in Sec. 3, it reveals a fundamental limitation: variance normalization alone cannot guarantee the balance between sufficiency (retaining predictive information) and minimality (removing redundancy). As a result, relying solely on variance-centric NOP ($\psi$) can produce representations with superior generalization and robustness, since higher IB values are empirically and theoretically linked to better generalization and robustness  (Kawaguchi et al., 2023; Alemi et al., 2016; Goldfeld & Polyanskiy, 2020; Saxe et al., 2019; Achille & Soatto, 2018).
@@
where $\{T_{l}\}_{l=1}^{L}\in\mathcal{T}_{1}\times\cdots\times\mathcal{T}_{L}\equiv\mathcal{T}^{\otimes L}$, $T_{0}=X$ and $\beta>0$ controls the trade-off between sufficiency and compression. Prior works (Kawaguchi et al., 2023; Saxe et al., 2019; Alemi et al., 2016; Goldfeld & Polyanskiy, 2020; Achille & Soatto, 2018) has shown that both generalization error and robustness positively correlate with the IB value in Eqn. (5), implying that aligning intermediate representations with the IB principle can improve generalization and robustness.
Building on this insight, we aim to enhance conventional variance-centric NOP ($\psi$) operator with an IB-inspired compression operation. Rather than treating normalization as a mere stabilizer, we reinterpret it as an information filter that produces representations that are both compact and predictive, bridging optimization stability and information-theoretic optimality. Directly optimizing Eqn.(5) is infeasible due to costly mutual information estimation over the unknown joint distribution $p_{X,Y}$, but it provides a guiding principle: normalization should go beyond variance standardization to encourage information-preserving activations. We introduce our IB-inspired normalization in Sec.4.2 and theoretically justify its superiority and improved generalization in Sec. 4.3.
(a) Under Gaussian (0,1) input.
(b) Gaussian (0,2) input.
Figure 1: Comparison of kernel density estimation for Gaussian inputs (mean 0, varying variance) under different compression operations: Standardization, IBNorm-L, IBNorm-T, and IBNorm-S ($\lambda=4$). Compression operations in IBNorm compress the tail of activations while adjusting higher-order statistics. IBNorm-L, IBNorm-T, and IBNorm-S, with the same $\lambda$, exhibit different abilities to compress the tails of activations. See more examples in Appendix C.3.

### v2

@@ Abstract
Normalization is fundamental to deep learning, but existing approaches such as BatchNorm, LayerNorm, and RMSNorm are variance-centric by enforcing zero mean and unit variance, stabilizing training without controlling how representations capture task-relevant information. We propose IB-Inspired Normalization (IBNorm), a simple yet powerful family of methods grounded in the Information Bottleneck principle. IBNorm introduces bounded compression operations that encourage embeddings to preserve predictive information while suppressing nuisance variability, yielding more informative representations while retaining the stability and compatibility of standard normalization. Theoretically, we prove that IBNorm achieves a higher IB value and tighter generalization bounds than variance-centric methods. Empirically, IBNorm outperforms BatchNorm, LayerNorm, and RMSNorm across large-scale language models (like LLaMA, GPT-2) and vision models (e.g., ResNet, ViT), with mutual information analysis confirming superior information bottleneck behavior. Code will be released publicly and is available in the supplementary now.
Machine Learning, ICML
@@ 1 Introduction
Normalization has become a cornerstone of deep learning, credited for stabilizing and accelerating training across domains. Techniques such as Batch Normalization (BN) (Ioffe and Szegedy, 2015), Layer Normalization (LN) (Ba et al., 2016), and RMSNorm (Zhang and Sennrich, 2019) are now standard in architectures spanning vision and language (Vaswani et al., 2017; Dosovitskiy et al., 2020; He et al., 2016b), underscoring normalization as a key driver of training stability and acceleration.
@@
Despite this success, existing normalization methods share a fundamental limitation: they are inherently variance-centric. By enforcing zero mean and unit variance, followed by rescaling and shifting, they improve the conditioning of the optimization problem (Huang et al., 2023, 2018; He et al., 2016a; Desjardins et al., 2015). However, this process operates purely on first- and second-order statistics, without offering any guidance on (intermediate) representation learning. Indeed, two representations may share identical mean and variance, but can encode drastically different amounts of task-relevant information (Raginsky et al., 2016; Achille and Soatto, 2018b) This raises a key question: can normalization be re-designed not only to stabilize training but also to shape representations toward sufficiency and generalization?
Prior works have explored this direction. For instance, NormalNorm (Eftekhari and Papyan, 2025) encourages Gaussian-like features via power transforms and additive noise, arguing that maximizing feature entropy improves representation learning. We argue that this objective is misaligned with generalization: Gaussian-like activations indiscriminately retain task-irrelevant nuisance factors and cannot filter task-irrelevant information (Kawaguchi et al., 2023; Tishby and Zaslavsky, 2015; Saxe et al., 2019). Additionally, NormalNorm destabilizes the zero-mean, unit-variance properties essential for efficient optimization (see Appendix B).
@@
Contributions. In this work, we move beyond variance-centric normalization and introduce IBNorm, a simple yet powerful normalization method grounded by the Information Bottleneck (IB) principle (Shamir et al., 2010; Tishby and Zaslavsky, 2015; Shwartz-Ziv and Tishby, 2017). Moreover, we also theoretically justify the superiority of IBNorm and its better generalization performance.
@@
First, we introduce an IB perspective to normalization, and propose IBNorm for generalization improvement. The IB principle seeks representations that preserve task-relevant information while discarding nuisances, thereby connecting normalization to generalization. Within this framework, we decompose existing normalization methods into three steps: (i) grouping features (e.g., across batch, channel, or dimension), (ii) standardization, and (iii) recovery via re-scaling and shifting. Our analysis reveals that the normalization operation itself is the key bottleneck governing information flow. Building on this insight, IBNorm introduces several different compression operators that selectively compress activations toward their means, providing explicit control over sufficiency-redundancy tradeoffs while remaining drop-in compatible with existing architectures.
@@
Second, we provide theoretical guarantees. Under the IB framework, we prove that IBNorm achieves a strictly larger IB value than variance-centric methods like LN and BN, thereby better balancing predictive sufficiency with nuisance suppression. Moreover, we establish a provably tighter generalization bound, explaining why IBNorm outperforms standard normalization in practice.
@@
Extensive experiments show the effectiveness of IBNorm across architectures and modalities. In language modeling, integrating IBNorm into LLaMA (Touvron et al., 2023b) and GPT-2 (Radford et al., 2019) outperforms LN, RMSNorm, and NormalNorm on LLM Leaderboard I/II by up to 8.75%. In vision, applying IBNorm to ResNet (He et al., 2016b) and ViT (Dosovitskiy et al., 2020) yields substantial gains, e.g., 3.98% improvement (ResNet-50) and 8.17% improvement (ViT) on ImageNet (Deng et al., 2009).
@@ 2.1 Normalization
Normalization plays a central role in modern deep learning, enabling faster convergence and improved generalization across natural language processing (NLP) (Vaswani et al., 2017; Touvron et al., 2023a, b; Dubey et al., 2024; Bai et al., 2023; Liu et al., 2024) and computer vision (CV) (He et al., 2016b; Dosovitskiy et al., 2020; Liu et al., 2022).
@@
Batch Normalization (BN) (Ioffe and Szegedy, 2015) normalizes activations across mini-batches to stabilize training, with numerous extensions such as Mean-only BN (Salimans and Kingma, 2016), $L^{p}$-Norm BN (Liao et al., 2016), Conditional BN (De Vries et al., 2017), and Whitening BN (Siarohin et al., 2018; Huang et al., 2018, 2019). Layer Normalization (LN) (Ba et al., 2016) was proposed for sequential and small-batch settings and has inspired variants including Dynamic LN (Kim et al., 2017), RMSNorm (Zhang and Sennrich, 2019), and Adaptive LN (Xu et al., 2019), which are now standard in Transformers (Vaswani et al., 2017; Dosovitskiy et al., 2020). Group Normalization (GN) (Wu and He, 2018) and Batch Group Normalization (BGN) (Summers and Dinneen, 2019) extend this idea by combining grouping across features, channels, or batches. Collectively, these methods underscore normalization as a structural tool for scaling depth and stabilizing optimization (Huang et al., 2023; Xu et al., 2019; Joudaki et al., 2023).
@@
Beyond variance-centric approaches, some works attempt to explicitly shape representation distributions. For example, Eftekhari and Papyan (2025) encourage Gaussian-like features via power transforms with additive noise, arguing that Gaussianity enhances representational capacity while noise acts as implicit regularization. We identify key theoretical and practical limitations in this approach. The method lacks a principled information-theoretic analysis within the context of deep neural networks. Specifically, the design of NormalNorm is based on the mutual information game (Theorem 2.1 (Eftekhari and Papyan, 2025)) which relies on strong assumptions regarding the first and second moments of the input and noise and neglects higher-order statistics. Consequently, the objective is misaligned with generalization: by prioritizing entropy maximization, the method indiscriminately preserves information, failing to filter task-irrelevant nuisance factors and leading to suboptimal performance. Practically, NormalNorm incurs computational overhead in estimating power parameters and introduces stochasticity that can destabilize training (see Appendix B and C.8).
@@
Broadly, most normalization methods rely on re-centering and rescaling to mitigate internal covariate shift, yet they overlook the information-theoretic properties of representations, such as sufficiency and redundancy. Crucially, even distribution-shaping methods like NormalNorm fail to explicitly balance preserving predictive information against suppressing nuisance factors. These gaps motivates our work. We propose IBNorm, which augments conventional normalization with compression operation designed to regulate information flow, guiding activations toward IB-optimal representations that enhance generalization and robustness.
2.2 Representation Learning with Explicit Information Bottleneck Objectives
The Information Bottleneck (IB) principle (Tishby et al., 2000; Shamir et al., 2010; Tishby and Zaslavsky, 2015) formulates representation learning as a trade-off between sufficiency and compression. Prior works (Tishby and Zaslavsky, 2015; Alemi et al., 2016; Kolchinsky et al., 2019; Belghazi et al., 2018; Hu et al., 2024) have explored this principle primarily through empirical analysis and leveraging explicit IB objectives during deep learning model training.
Methods such as deep variational information bottleneck (VIB) (Alemi et al., 2016), nonlinear information bottleneck (Kolchinsky et al., 2019), and mutual information neural estimation (MINE) (Belghazi et al., 2018) leverage the IB principle by explicitly estimating mutual information through auxiliary neural MI estimators and incorporating these estimates as IB-regularization terms during training. These approaches rely on encoder-decoder architectures, require large datasets to train accurate MI estimators prior to training the primary model, and introduce additional MI-based loss functions that cause both computational and hyperparameter overhead. The reliance on auxiliary estimators also imposes architectural constraints and increases the computational burden, a limitation that becomes pronounced in autoregressive LLMs. Furthermore, as demonstrated by MINE, accurate MI estimation requires substantial sample complexity, making these methods expensive and often unstable when applied to large-scale models. Recent techniques such as structured probabilistic coding (Hu et al., 2024) incorporate IB-inspired coding mechanisms but are confined to specific encoder architectures and inherit similar computational and optimization challenges.
Representation learning methods based on explicit IB objectives face many limitations: they impose heavy computational overhead due to auxiliary MI-estimator networks, introduce architectural constraints, and require large sample sizes for accurate mutual information estimation. Detailed discussion can be found in Appendix A.1 and A.2.
To effectively integrate the IB principle into modern deep learning models, IBNorm introduces the first framework to internalize the IB principle directly within the normalization operation. By redesigning the normalization layer to act as an information filter, IBNorm achieves the benefits of IB-guided compression—improved generalization and robustness—without computationally expensive auxiliary estimators, external losses, or architectural constraints.
@@ 3 Preliminary
Here we introduce information bottleneck (IB) framework (Shamir et al., 2010; Tishby and Zaslavsky, 2015) which forms the basis of our approach in Sec. 4.
Consider two random variables $X$ and $Y$ with joint probability densities $p(x,y)$ over the space $\mathcal{X}\times\mathcal{Y}$. We can define the mutual information between $X$ and $Y$:
$\vskip-4.49997ptI(X;Y)=\int_{x\in\mathcal{X}}\int_{y\in\mathcal{Y}}p(x,y)\,\log\frac{p(x,y)}{p(x)\,p(y)}\,dx\,dy.$ (1)
Mutual information measures dependencies between random variables, and can be understood as how much knowing $X$ reduces the uncertainty in $Y$ or vice versa (Hjelm et al., 2018; Gabrié et al., 2018; Kong et al., 2019).
Building on this, the IB framework formulates representation learning as a trade-off between two goals: 1) minimality which compresses $X$ by reducing $I(X;T)$; and 2) sufficiency which preserves information about $Y$ by maximizing $I(Y;T)$. This is captured by the IB objective:
$\max_{T}\Big[I(Y;T)-\beta I(X;T)\Big],$ (2)
where $\beta>0$ (Tishby and Zaslavsky, 2015). Here, $T$ must satisfy the Markov chain TXYT-X-Y, and the optimal representation $T^{*}$ is given by the conditional distribution $p_{T^{*}|X}(t|x)$ with marginal $p_{T^{*}}(t)=\int p_{T^{*}|X}(t|x)p_{X}(x)dx$ governed by self-consistency equations (Tishby et al., 2000).
Intuitively, maximizing $I(Y;T)$ ensures that the representation $T$ preserving information relevant for predicting the target $Y$, thereby enhancing predictive power and interpretability. Concurrently, penalizing $I(X;T)$ enforces compression, which eliminates irrelevant variability in the representation and, in turn, enhances generalization and robustness of the representation (Kawaguchi et al., 2023; Goldfeld and Polyanskiy, 2020; Saxe et al., 2019; Federici et al., 2020).
@@ 4 Methodology
Inspired by the IB principle introduced in Sec. 3, we present how it motivates an effective normalization strategy — an essential component of modern deep learning architectures.
@@ 4.1 Normalization Design via IB Principle
Consider a data space $\mathcal{X}$ and label space $\mathcal{Y}$ with a fixed joint distribution $\mathbb{P}(X,Y)\in\mathcal{P}(\mathcal{X}\times\mathcal{Y})$. A feedforward neural network $f(\cdot;\theta)$, parameterized by $\theta$, maps an input $X\in\mathcal{X}$ to a prediction $\hat{Y}=f(X;\theta)$, aiming to approximate the ground-truth label $Y$. Here, $f(\cdot;\theta)$ can represent various architectures, such as convolutional neural networks (e.g., ResNet (He et al., 2016b)) and transformer-based networks (e.g., large language models like (Touvron et al., 2023b) or vision transformer (Dosovitskiy et al., 2020)).
For an $L$-layer network with normalization, we write the intermediate representation after the $l$-th normalization as:
$T_{l}=\Phi_{l}\circ h_{l}\circ\cdots\circ\Phi_{1}\circ h_{1}(X)\in\mathcal{T}_{l},\quad l\in\{1,\cdots,L\},$ (3)
where $\mathcal{T}_{l}$ is a measurable space, $h_{i}$ is a transformation block with parameter $\theta_{i}$, and $\Phi_{i}$ is the normalization layer, e.g., layer normalization. We use symbol $\circ$ to represent the composition of functions. Following (Huang et al., 2023), any normalization layer $\Phi$ can be decomposed as:
@@
This decomposition unifies existing normalization methods. For instance, BN (Ioffe and Szegedy, 2015), LN (Ba et al., 2016), and GN (Wu and He, 2018) differ mainly in their choice of NAP ($\zeta$): BN normalizes across the batch dimension, LN across the feature dimension, and GN across groups of features. In all cases, the NOP enforces zero-mean and unit-variance statistics, while the NRR rescales and shifts the result. Crucially, among them, NAP ($\zeta$) is often dictated by the network architecture or task domain (e.g., batch-based normalization favored by CNNs, feature-based normalization for Transformers). Likewise, NRR ($\eta$) is linear and invertible, and therefore only re-parameterizes activations without altering their information content. This leaves the NOP ($\psi$) as the main degree of freedom: it governs how information is filtered and thus determines whether normalization merely stabilizes optimization or also enhances the quality of learned representations.
However, most existing NOP ($\psi$) operations are fundamentally variance-centric, enforcing zero-mean and unit-variance scaling of activations. While stabilizing and accelerating training, it overlooks whether the resulting representations retain task-relevant information. From the IB perspective in Sec. 3, it reveals a fundamental limitation: variance normalization alone cannot guarantee the balance between sufficiency (retaining predictive information) and minimality (removing redundancy). As a result, relying solely on variance-centric NOP ($\psi$) can produce representations with superior generalization and robustness, since higher IB values are empirically and theoretically linked to better generalization and robustness  (Kawaguchi et al., 2023; Alemi et al., 2016; Goldfeld and Polyanskiy, 2020; Saxe et al., 2019; Achille and Soatto, 2018a).
@@
where $\{T_{l}\}_{l=1}^{L}\in\mathcal{T}_{1}\times\cdots\times\mathcal{T}_{L}\equiv\mathcal{T}^{\otimes L}$, $T_{0}=X$ and $\beta>0$ controls the trade-off between sufficiency and compression. Prior works (Kawaguchi et al., 2023; Saxe et al., 2019; Alemi et al., 2016; Goldfeld and Polyanskiy, 2020; Achille and Soatto, 2018a) has shown that both generalization error and robustness positively correlate with the IB value in Eqn. (5), implying that aligning intermediate representations with the IB principle can improve generalization and robustness.
Building on this insight, we aim to enhance conventional variance-centric NOP ($\psi$) operator with an IB-inspired compression operation. Rather than treating normalization as a mere stabilizer, we reinterpret it as an information filter that produces representations that are both compact and predictive, bridging optimization stability and information-theoretic optimality. Directly optimizing Eqn. (5) as an additional IB-based training loss is infeasible due to two main reasons: (1) the computational cost of estimating mutual information using auxiliary MI neural estimators over the unknown joint distribution $p_{X,Y}$, and (2) the significant training overhead and sample complexity associated with such MI neural estimators (see details in Appendix A.2). However, it provides a guiding principle: normalization should go beyond variance standardization to encourage information-preserving activations. We introduce our IB-inspired normalization in Sec. 4.2 and theoretically justify its superiority and improved generalization in Sec. 4.3.
Figure 1: Comparison of kernel density estimation for Gaussian inputs (mean 0, varying variance) under different compression operations: Standardization, IBNorm-L, IBNorm-T, and IBNorm-S ($\lambda=4$). See more examples in Appendix C.6.